# CaVE: A Cone-Aligned Approach for Fast Predict-then-optimize with Binary Linear Programs

### 12 Dec 2023 v1

@@ Abstract
The end-to-end predict-then-optimize framework, also known as decision-focused learning, has gained popularity for its ability to integrate optimization into the training procedure of machine learning models that predict the unknown cost (objective function) coefficients of optimization problems from contextual instance information. Naturally, most of the problems of interest in this space can be cast as integer linear programs. In this work, we focus on binary linear programs (BLPs) and propose a new end-to-end training method for predict-then-optimize. Our method, Cone-aligned Vector Estimation (CaVE), aligns the predicted cost vectors with the cone corresponding to the true optimal solution of a training instance. When the predicted cost vector lies inside the cone, the optimal solution to the linear relaxation of the binary problem is optimal w.r.t. to the true cost vector. Not only does this alignment produce decision-aware learning models, but it also dramatically reduces training time as it circumvents the need to solve BLPs to compute a loss function with its gradients. Experiments across multiple datasets show that our method exhibits a favorable trade-off between training time and solution quality, particularly with large-scale optimization problems such as vehicle routing, a hard BLP that has yet to benefit from predict-then-optimize methods in the literature due to its difficulty.
@@ 1 Introduction
Theoretical and experimental results reported over the past few years, starting with Elmachtoub and Grigas [7] and Ban and Rudin [2], have demonstrated the need for end-to-end training of Machine Learning (ML) models that predict the cost coefficients of optimization problems. This contrasts with the more traditional two-stage approach, where an ML model is first trained to minimize regression loss, and then its predictions are applied to new test instances. This conventional approach often leads to substantial regret in terms of the quality of the solutions obtained, especially when the training set is small. Given that such predict-then-optimize settings are commonly encountered in many applications (e.g., predicting product demand to manage inventory or travel time on a road network to route trucks), researchers in ML and optimization have proposed a wide range of end-to-end training methods, many of which have been surveyed and compared recently [28, 24, 16].
@@
With a few exceptions [6, 11], these methods largely follow the now prevalent mini-batch stochastic gradient descent training algorithm. This process involves the following steps in each iteration: (0) a small batch of training instances is fed into the ML model; (1) the model predicts the cost coefficients; (2) a loss function, incorporating the concept of “decision error” such as regret, is calculated; (3) the gradients of this loss w.r.t. the parameters of the ML model are computed using backpropagation; and (4) a gradient descent step is employed to update the model parameters. A common feature of many of these methods is the necessity of solving the optimization problem for each training instance at least once in the forward pass (Steps (1, 2)). Since these repeated calls to the (integer) optimization solver represent a significant computational bottleneck, there have been attempts to improve efficiency by replacing them with much cheaper linear optimization calls [17], solution caching [20, 14], and function approximation [27, 26, 9]. However, these measures often come at a sacrifice in solution quality, as we empirically demonstrate in this paper.
@@
In this work, we are interested in efficient end-to-end training of ML models that predict cost coefficients of challenging binary linear optimization problems, a prime example of which is the Capacitated Vehicle Routing Problem (CVRP) [29]. CVRP is defined on a graph comprising customers and a depot, where binary variables represent the edges of the graph and indicate whether or not a vehicle traverses that edge. The number of vehicles, their capacities, and customer demands are fixed and known. Linear constraints capture the requirements of valid tours for all vehicles starting and ending at the depot, adhering to vehicle capacity limits, and visiting each customer exactly once. In the predict-then-optimize setting, each CVRP instance is associated with a “feature vector” (e.g., weather conditions, time-of-day, whether it is a holiday or not, etc.) that is predictive of the (unknown) travel times on the graph’s edges, which are also the cost coefficients of the objective function, namely the total travel time. CVRP is notoriously hard to solve, even for tens of customers, making end-to-end training of ML models extremely time-consuming.
@@
A distinguishing feature of our CaVE loss functions is that they do not require solving the original optimization problem during training. Instead, they rely on easier projection problems, continuous and quadratic, which are significantly faster to solve. Our key insight is as follows: By ensuring that the predicted cost vector falls inside a specific cone, namely one that corresponds to the optimal solution under the true cost vector, we are likely to recover this optimal solution. The binding (or active) constraints at the optimal solution define this critical cone. To align the ML model’s predicted cost vector with the cone, we minimize the angle between the prediction and the cone; this is done through projection onto the cone, our main optimization routine. We show that CaVE trains ML models in a fraction of the time required by state-of-the-art methods such as SPO+ [7] and PFYL [3] while yielding equally effective cost predictions as measured by regret on unseen test instances.
@@ 2 Related Work
A notable advancement in this area is the KKT-based method: Amos and Kolter [1] obtain both optimal solutions and gradients by solving a linear system derived from the KKT conditions. Wilder et al. [31] adapted the method for linear programs (LPs) by incorporating a small quadratic term into the objective function, while Mandi and Guns [15] introduced a logarithmic barrier term. Furthermore, Ferber et al. [8] employ the cutting-plane method, which allows for integer variables. These KKT-based implicit differentiation methods require the use of specialized solving algorithms, which inherently limit their flexibility and often compromise the efficiency of solving problems such as linear programming. Furthermore, aside from the time-intensive cutting-plane method, KKT-based approaches generally struggle to tackle discrete models.
@@
In contrast, other methodologies approach the optimization solver as a black box, functioning independently of the solver and algorithm. Elmachtoub and Grigas [7] propose a convex surrogate of regret for linear objective functions, which has nonzero subgradients. In contrast to this convex and theoretically sound loss function, Pogančić et al. [23] propose to linearly interpolate the loss function, transforming it into a piecewise linear function. As a specific case of the interpolation method, Sahoo et al. [25] adopt a more direct strategy by using the negative identity matrix as the surrogate optimization gradient. Niepert et al. [21] extend the interpolated method with perturbations with random noise. In the context of linear objective functions, the perturbation approach [3, 4] involves adding a perturbation to the predicted cost coefficients to smooth the optimization function and further constructs a loss function based on duality.
Since the training process involves solving the optimization problem in each iteration, many solutions naturally accumulate as samples. Under the assumption that the feasible region remains fixed, these solutions are all feasible. Therefore, Mulamba et al. [20] proposes a contrastive loss designed to maximize the distinction between suboptimal solutions in the sample and the optimal solution. In addition, it utilizes these accumulated solutions as a cache, effectively reducing computational cost. Inspired by the contrastive approach, Mandi et al. [14] employed “learning-to-rank” [13] by ranking the objective value of the cached solutions as well as the optimal one.
@@
Given the computational inefficiency often associated with solving optimization, it is attractive to avoid it, such as function approximation. The critical component of function approximation is a learnable surrogate function, which is learned to mimic the original objective or loss function. Shah et al. [27, 26] samples datasets and employs an additional neural network model. This model is trained to approximate the actual loss of a decision. In doing so, they effectively deploy this approximate loss in end-to-end training. In addition, Ferber et al. [9] take a unique approach by substituting the nonlinear objective function with a linear surrogate. With the approximation method, the complexity of directly dealing with the original function is significantly reduced, allowing for more efficient learning and optimization. However, the accuracy of the approximate loss function directly impacts the final model performance. In practice, training a model to learn and approximate a specific function effectively can be a challenging endeavour.
@@ 3.1 Definitions and Notation
For the sake of clarity, we can define a binary linear program as follows: There are $d$ binary decision variables $\bm{w}$, each taking a value of either 00 or $1$. The cost coefficients associated with these decision variables are represented by $\bm{c}\in\mathbb{R}^{d}$; the constraints are $\bm{A}\bm{w}\leq\bm{b}$, where $\bm{A}\in\mathbb{R}^{k\times d}$ and $\bm{b}\in\mathbb{R}^{k}$:
@@
$\bm{w}\in\{\bm{0},\bm{1}\}.$
Let $S$ represent the feasible region of the problem; then the optimal solution is expressed as $\bm{w}^{*}(\bm{c})=\operatorname*{arg\,min}_{\bm{w}\in S}\bm{c}^{\intercal}\bm{w}$. Given its computational complexity, the process of determining $\bm{w}^{*}(\bm{c})$ can be extremely time-consuming, especially when the number of decision variables and constraints is large.
In the predict-then-optimize setting, the coefficients $\bm{c}$ are unknown and associated with a feature vector $\bm{x}\in\mathbb{R}^{p}$. This relationship facilitates the use of an ML model $g(\bm{x},\bm{\theta})$ to estimate the predicted coefficients $\hat{\bm{c}}$. In this model, $\bm{\theta}$ represents the learnable parameters, which are adjusted to minimize a decision loss $\mathcal{L}(\cdot)$, a metric that quantifies the discrepancy between the true optimal solution $\bm{w}^{*}(\bm{c})$ and the solution derived from prediction $\hat{\bm{c}}$.
@@ 3.2 Metric
To measure the quality of the decision, the concept of regret is introduced. Regret is defined as the optimality gap between the optimal solution $\bm{w}^{*}(\hat{\bm{c}})$ obtained using the predicted coefficients $\hat{\bm{c}}$, and the true optimal solution $\bm{w}^{*}(\bm{c})$ obtained using the actual coefficients $\bm{c}$. It is expressed in the following equation:
$\mathcal{L}_{\text{Regret}}(\hat{\bm{c}},\bm{c})=\bm{c}^{\intercal}\big{(}\bm{w}^{*}\left(\hat{\bm{c}}\right)-\bm{w}^{*}\left(\bm{c}\right)\big{)}.$ (2)
In line with Elmachtoub and Grigas [7], we adopt normalized regret, which serves as an adjusted metric that takes into account the scale of the problem, providing a more standardized and comparable measure:
$\frac{\sum_{i=1}^{n}{\mathcal{L}_{\text{Regret}}\left(\hat{\bm{c}}_{i},\bm{c}_{i}\right)}}{\sum_{i=1}^{n}{\left|\bm{c}^{\intercal}\bm{w}^{*}(\bm{c})\right|}}.$ (3)
The regret is best understood as follows: if a method records a test regret of $0.07$ for example, it means that it produces solutions that are $7\%$ worse than the true optimal solutions under the true by unknown cost vectors.
@@ 4.1 Optimal Cones and Subcones
Figure 1: Illustration of an optimal cone: The cost vectors $\bm{c}_{1}$ and $\bm{c}_{2}$ produce the same optimal solution if and only if within this cone.
@@
For a (continuous) LP, one can associate an optimal cone with a given cost vector $\bm{c}\in\mathbb{R}^{d}$. Within this cone, all cost coefficients yield the same optimal solution $\bm{w}^{*}(\bm{c})$. As depicted on the right of Fig. 1, the optimal cone is constructed from the conical combination of the binding constraints at $\bm{w}^{*}(\bm{c})$: Each vector in the cone can be expressed as a non-negative combination of the coefficient vectors to the binding constraints. Now consider the case of a binary linear program. If linear constraints correspond to the convex hull of integer points of the BLP, the same logic to obtain the optimal cone of an integer solution would have been applied. The left side of Fig. 2 illustrates such an optimal cone for a BLP. The cone is defined by the binding constraints $w_{2}\leq 1$ (left vector) and $w_{1}+w_{2}\leq 1$ (right vector). However, we typically do not operate on the convex hull of a BLP, necessitating an alternate definition for the cones that we will attempt to align the predictive model’s predictions with.
To this end, we introduce the concept of an optimal subcone. For a BLP with optimal solution $\bm{w}^{*}(\bm{c})$, the optimal subcone is the optimal cone of the same solution but with the BLP’s LP relaxation instead. In Fig. 2, the cone on the left side is the actual optimal cone for the BLP; the cone on the right side is the optimal subcone. All rays within the subcone are also part of the optimal cone. Although the LP relaxation leads to an expanded feasible region, all binary feasible solutions of the original BLP remain feasible vertices in the LP. As such, recovering an optimal subcone for a given solution $\bm{w}^{*}(\bm{c})$ is trivial: For a BLP problem defined as in Equation 1, its optimal subcone is the set of vectors that can be expressed as nonnegative combinations of the coefficient vectors of binding constraints $\widetilde{\bm{A}}$, i.e., constraints that satisfy $\bm{A}_{j}\bm{w}^{*}(\bm{c})=b_{j}$, alongside the variable bounds for which we have either $w_{i}^{*}(\bm{c})=0$ or $w_{i}^{*}(\bm{c})=1$.

### 15 Mar 2024 v2

@@ Abstract
The end-to-end predict-then-optimize framework, also known as decision-focused learning, has gained popularity for its ability to integrate optimization into the training procedure of machine learning models that predict the unknown cost (objective function) coefficients of optimization problems from contextual instance information. Naturally, most of the problems of interest in this space can be cast as integer linear programs. In this work, we focus on binary linear programs (BLPs) and propose a new end-to-end training method to predict-then-optimize. Our method, Cone-aligned Vector Estimation (CaVE), aligns the predicted cost vectors with the normal cone corresponding to the true optimal solution of a training instance. When the predicted cost vector lies inside the cone, the optimal solution to the linear relaxation of the binary problem is optimal. This alignment not only produces decision-aware learning models, but also dramatically reduces training time as it circumvents the need to solve BLPs to compute a loss function with its gradients. Experiments across multiple datasets show that our method exhibits a favorable trade-off between training time and solution quality, particularly with large-scale optimization problems such as vehicle routing, a hard BLP that has yet to benefit from predict-then-optimize methods in the literature due to its difficulty.
@@ 1 Introduction
Theoretical and experimental results reported over the past few years, starting with Elmachtoub and Grigas [7] and Ban and Rudin [2], have demonstrated the need for end-to-end training of Machine Learning (ML) models that predict the cost coefficients of optimization problems. This contrasts with the more traditional two-stage approach, where an ML model is first trained to minimize regression loss for prediction, and then its predictions are applied to new test instances for decision. This conventional approach often leads to substantial regret in terms of the quality of the solutions obtained, especially when the training set is small. Given that such predict-then-optimize settings are commonly encountered in many applications (e.g., predicting product demand to manage inventory or travel time on a road network to route trucks), researchers in ML and optimization have proposed a wide range of end-to-end training methods, many of which have been recently surveyed and compared [28, 24, 16].
@@
With a few exceptions [6, 11], these methods largely follow the now prevalent mini-batch stochastic gradient descent training algorithm. This process involves the following steps in each iteration: (0) a small batch of training instances is fed into the ML model; (1) the model predicts the cost coefficients; (2) a loss function, incorporating the concept of “decision error” such as regret, is calculated; (3) the gradients of this loss w.r.t. the parameters of the ML model are computed using backpropagation; and (4) a gradient descent step is employed to update the model parameters. A common feature of many of these methods is the necessity to solve the optimization problem for each training instance at least once in the forward pass (Steps (1, 2)). Since these repeated calls to the (integer) optimization solver represent a significant computational bottleneck, there have been attempts to improve efficiency by replacing them with much cheaper linear optimization calls [17], solution caching [20, 14], and function approximation [27, 26, 9]. However, these measures often come at a sacrifice in solution quality, as we empirically demonstrate in this paper.
@@
In this work, we are interested in efficient end-to-end training of ML models that predict cost coefficients of challenging binary linear optimization problems, a prime example of which is the Capacitated Vehicle Routing Problem (CVRP) [29]. CVRP is defined on a graph comprising customers and a depot, where binary variables represent the edges of the graph and indicate whether or not a vehicle traverses that edge. The number of vehicles, their capacities, and customer demands are fixed and known. Linear constraints capture the requirements of valid tours for all vehicles that start and end at the depot, comply with vehicle capacity limits, and visit each customer exactly once. In the predict-then-optimize setting, each CVRP instance is associated with a “feature vector” (e.g., weather conditions, time-of-day, whether it is a holiday or not, etc.) that is predictive of the (unknown) travel times on the graph’s edges, which are also the cost coefficients of the objective function, namely the total travel time. CVRP is notoriously hard to solve, even for tens of customers, making end-to-end training of ML models extremely time-consuming.
@@
A distinguishing feature of our CaVE loss functions is that they do not require solving the original optimization problem during training. Instead, they rely on easier projection problems, continuous and quadratic, which are significantly faster to solve. Our key insight is as follows: By ensuring that the predicted cost vector falls inside a specific cone, namely, one that corresponds to the optimal solution under the true cost vector, we are able to recover this optimal solution. The binding (or active) constraints at the optimal solution define this critical cone. To align the ML model’s predicted cost vector with the cone, we need to minimize the angle between the prediction and the cone; this is done through projection onto the cone, our main optimization routine. We show that CaVE trains ML models in a fraction of the time required by state-of-the-art methods such as SPO+ [7] and PFYL [3] while yielding equally effective cost predictions as measured by regret in unseen test instances.
@@ 2 Related Work
A notable advancement in this area is the KKT-based method: Amos and Kolter [1] obtain both optimal solutions and gradients by solving a linear system derived from KKT conditions. Wilder et al. [31] adapted the method for linear programs (LPs) by incorporating a small quadratic term into the objective function, while Mandi and Guns [15] introduced a logarithmic barrier term. Furthermore, Ferber et al. [8] employ the cutting-plane method, which allows for integer variables. These KKT-based implicit differentiation methods require the use of specialized solving algorithms, which inherently limit their flexibility and often compromise the efficiency of solving problems such as linear programming. Furthermore, aside from the time-intensive cutting-plane method, KKT-based approaches generally struggle to tackle discrete models.
@@
In contrast, other methodologies approach the optimization solver as a black box, functioning independently of the solver and algorithm. Elmachtoub and Grigas [7] propose a convex surrogate of regret for linear objective functions, which has nonzero subgradients. In contrast to this convex and theoretically sound loss function, Pogančić et al. [23] present a linear interpolation, transforming optimization into a piecewise linear function. As a specific case of the interpolation method, Sahoo et al. [25] adopt a more direct straight-through estimator by using the negative identity matrix as the surrogate optimization gradient. Niepert et al. [21] extend the interpolated method with perturbations with random noise. In the context of linear objective functions, more perturbation approaches [3, 4] involve adding a perturbation to the predicted cost coefficients to smooth the optimization function and further construct a loss function based on duality.
Since the training process involves solving the optimization problem in each iteration, many solutions naturally accumulate as samples. Under the assumption that the feasible region remains fixed, these solutions are all feasible. Therefore, Mulamba et al. [20] proposes a contrastive loss designed to maximize the distinction between suboptimal solutions in the sample and the optimal solution. Additionally, it utilizes these accumulated solutions as a cache, effectively reducing computational cost. Inspired by the contrastive approach, Mandi et al. [14] employed “learning-to-rank” [13] by ranking the objective value of the cached solutions.
@@
Due to the computational inefficiency often encountered in solving optimization problems, function approximation methods that do not require constraint optimization are appealing. The critical component of function approximation is a learnable surrogate function, which is learned to mimic the original objective or loss function. Shah et al. [27, 26] samples datasets and employs an additional neural network model. This model is trained to approximate the actual loss of a decision. In doing so, they effectively deploy this approximate loss in end-to-end training. With the approximation method, the complexity of directly dealing with the original function is significantly reduced, allowing for more efficient learning. However, the accuracy of the approximate loss function directly impacts the final model performance. In practice, training a model to learn and approximate a specific function effectively can be a challenging endeavour.
@@ 3.1 Definitions and Notation
For the sake of clarity, we can define a binary linear program as follows: There are binary decision variables $\bm{w}\in\{\bm{0},\bm{1}\}^{d}$. The cost coefficients associated with these decision variables are represented by $\bm{c}\in\mathbb{R}^{d}$; the constraints are $\bm{A}\bm{w}\leq\bm{b}$, where $\bm{A}\in\mathbb{R}^{k\times d}$ and $\bm{b}\in\mathbb{R}^{k}$:
@@
$\bm{w}\in\{\bm{0},\bm{1}\}^{d}.$
Let $\bm{\Omega}$ represent the feasible region of the problem; then the optimal solution is expressed as $\bm{w}^{*}(\bm{c})=\operatorname*{arg\,min}_{\bm{w}\in\bm{\Omega}}\bm{c}^{\intercal}\bm{w}$. Given its computational complexity, the process of determining $\bm{w}^{*}(\bm{c})$ can be extremely time-consuming, especially when the number of decision variables and constraints is large.
In the predict-then-optimize setting, the coefficients $\bm{c}$ are unknown and linked to a feature vector $\bm{x}\in\mathbb{R}^{p}$. This relationship facilitates the use of an ML model $g(\bm{x},\bm{\theta})$ to estimate the predicted coefficients $\hat{\bm{c}}$. In this model, $\bm{\theta}$ represents the learnable parameters, which are adjusted to minimize a decision loss $\mathcal{L}(\cdot)$, a metric that quantifies the discrepancy between the true optimal solution $\bm{w}^{*}(\bm{c})$ and the solution derived from the prediction $\hat{\bm{c}}$.
@@ 3.2 Metric
To measure the quality of the decision, regret is introduced by Elmachtoub and Grigas [7], which is defined as the absolute gap between the objective value of the solution $\bm{w}^{*}(\hat{\bm{c}})$ obtained using the predicted coefficients $\hat{\bm{c}}$, and the optimal value of the solution $\bm{w}^{*}(\bm{c})$ obtained using the actual coefficients $\bm{c}$. It is expressed in the following equation:
$\mathcal{L}_{\text{Regret}}(\hat{\bm{c}},\bm{c})=\bm{c}^{\intercal}\big{(}\bm{w}^{*}(\hat{\bm{c}})-\bm{w}^{*}(\bm{c})\big{)}.$ (2)
In line with Elmachtoub and Grigas [7], we adopt normalized regret, which serves as an adjusted metric that takes into account the scale of the problem, providing a more standardized and comparable measure with $n$ samples:
$\frac{\sum_{i=1}^{n}{\mathcal{L}_{\text{Regret}}\left(\hat{\bm{c}}_{i},\bm{c}_{i}\right)}}{\sum_{i=1}^{n}{\left|\bm{c}_{i}^{\intercal}\bm{w}_{i}^{*}(\bm{c}_{i})\right|}}.$ (3)
The regret is best understood as follows: If a method records a test regret of $0.07$ for example, it means that it produces solutions that are $7\%$ worse than the true optimal solutions under the true but unknown cost vectors.
@@ 4.1 Optimal Cones and Subcones
Figure 1: Illustration of a normal cone: the cost vectors $\bm{c}_{1}$ and $\bm{c}_{2}$ produce the same optimal solution if and only if they lie within this cone.
@@
In (continuous) LP, one can associate a normal cone as optimal cone with a given cost vector $\bm{c}\in\mathbb{R}^{d}$. Within this cone, all cost coefficients yield the same optimal solution $\bm{w}^{*}(\bm{c})$. As depicted on the right of Fig. 1, the construction of the normal cone leverages the conical combination of the binding constraints $\widetilde{\bm{A}}(\bm{c})$ at $\bm{w}^{*}(\bm{c})$: each vector within the cone can be represented as a non-negative combination of the binding constraints coefficients $\bm{a}_{j}\in\widetilde{\bm{A}}(\bm{c})$, which can be written as
$\sum_{\bm{a}_{j}\in\widetilde{\bm{A}}(\bm{c})}{\lambda}_{i}\bm{a}_{j},\forall\bm{\lambda}\geq\bm{0},$
Now consider the case of a binary linear program. If linear cuts $\bm{A}^{\prime}(\bm{c})$ correspond to the convex hull of integer points of the BLP, the same logic to obtain the optimal cone of an integer solution would have been applied. This parallel is illustrated on the right side of Figure 2 for a BLP optimal cone, defined as
$\mathcal{C}^{*}(\bm{c})=\left\{\bm{v}\in\mathbb{R}^{d}:\bm{v}=\sum_{\bm{a}_{j}\in\bm{A}^{\prime}(\bm{c})}\lambda_{i}\bm{a}_{i},\forall\bm{\lambda}\geq\bm{0}\right\}.$
Here, the cone is delineated by the cuts $w_{1}\geq 0$, $w_{2}\leq 1$ and $w_{1}+w_{2}\leq 1$. However, we typically do not operate on the convex hull of a BLP, necessitating an alternate definition for the cones with which we will attempt to align the predictions.
To this end, we keep the normal cone of a BLP as the optimal subcone:
$\mathcal{SC}^{*}(\bm{c})=\left\{\bm{v}\in\mathbb{R}^{d}:\bm{v}=\sum_{\bm{a}_{j}\in\widetilde{\bm{A}}(\bm{c})}\lambda_{i}\bm{a}_{i},\forall\bm{\lambda}\geq\bm{0}\right\}.$
For a BLP with an optimal solution $\bm{w}^{*}(\bm{c})$, the optimal subcone is the optimal cone of the same solution, but with the LP relaxation of the BLP instead. In Fig. 2, the cone on the left side is the optimal cone for the BLP; the cone on the right side is the optimal subcone. Since $\mathcal{SC}^{*}(\bm{c})\subset\mathcal{C}^{*}(\bm{c})$, all rays within the subcone also belong to the optimal cone. Although the LP relaxation leads to an expanded feasible region, all binary feasible solutions of the original BLP remain feasible vertices in the LP. As such, recovering an optimal subcone for a given solution $\bm{w}^{*}(\bm{c})$ is trivial: for a BLP problem defined as in Equation 1, its optimal subcone is the set of vectors that can be expressed as nonnegative combinations of the coefficient vectors $\bm{a}_{j}\in\widetilde{\bm{A}}(\bm{c})$ that satisfy $\bm{a}_{j}\bm{w}^{*}(\bm{c})=b_{j}$, including the variable bounds for which we have either $w_{i}^{*}(\bm{c})=0$ or $w_{i}^{*}(\bm{c})=1$.